NvidiaIntely (nvidiaintely):

A 300 gram sample of pure water exerts a vapor pressure of 750 millimeters of mercury on the walls of its container. If 0.27 moles of glucose (C6H12O6) is added to the water, what will be the vapor pressure of the resulting solution? 0.984 mm Hg 12.1 mm Hg 738 mm Hg 748 mm Hg

NvidiaIntely (nvidiaintely):

@sshayer @sweetburger @mathstudent55

NvidiaIntely (nvidiaintely):

@AloneS

NvidiaIntely (nvidiaintely):

@Nnesha

OpenStudy (kg1975):

Psolution is the vapor pressure of the solution Χsolvent is mole fraction of the solvent P0solvent is the vapor pressure of the pure solvent mole fraction = moles of solute/total solution moles moles glucose = .27 finding moles water = 300/18 = 16.67 total moles solution = 16.67 + .27 = 16.93 mole fraction = .27/(16.93) = 0.0159 Rauolt's Law: Psolution = (Χsolvent)(P0solvent) = (0.0159)(1 atm) i think 1 atm = 750 mmHg Psolution = 0.984

OpenStudy (kg1975):

believe me or not :T like i care but a few people aint gonna stop me from helping no matter what u think

NvidiaIntely (nvidiaintely):

The answer is 738 mmHg.

OpenStudy (kg1975):

sure.. if u knew why did u ask even tho ur wrong

OpenStudy (quickstudent):

@kg1975 .............he simply submitted the quiz, and his teacher said that that was the real answer.